# Adaptive Uncertainty Quantification for Generative AI

### v1

@@ 6 Concluding Remarks
Conformal prediction serves to provide some statistical clarity in the otherwise uncharted waters of modern black-box deep learning. Conformal tree allows for local adaptation and conditional coverage guarantees based on a self-grouping procedure that is robust to the inclusion of an additional i.i.d test point. This can be a valuable tool for quantifying uncertainty in black-box learning applications where the model outputs a single point prediction using training data that has been obscured from the end user. Future work may include extending the framework beyond trees to other classes of calibration models.
@@ A.2 Proof of Lemma 1
For the first case, the only possibility is that new candidates are created if the exclusion of $(X_{n+1},S_{n+1})$ changes the number of datapoints inside a potential leaf from $m$ to $m-1$. For this to occur, it requires that exactly $m$ points are contained in the potential leaf node containing $X_{n+1}$. Denote by $\mathscr{X}_{k(X_{n+1})}$ the leaf of $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ containing $X_{n+1}$. We define the event such that removing $X_{n+1}$ doest not change the candidacy (i.e. eligibility for splitting) of the node $\mathscr{X}_{k(X_{n+1})}$ as
@@
$\delta(n,m):=\frac{2}{m}+\binom{n+1}{m}\left(\frac{m}{n+1}\right)^{m}\left(1-\frac{m}{n+1}\right)^{n+1-m}$ (A.9)
@@
To upper bound the probability of $\mathcal{E}^{c}_{\mathrm{cand}}$, we consider the worst-case distribution over $X_{1},\ldots,X_{n+1}$ by considering the worst-case amount of probability mass assigned to $\mathscr{X}_{k}$. That is,
$\mathbb{P}(\mathcal{E}_{\mathrm{cand}}\mid\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})=\widetilde{\mathfrak{X}},X_{n+1}\in\mathscr{X}_{k})$ $\leq\max_{p\in[0,1]}\binom{n+1}{m}p^{m}(1-p)^{n+1-m}$
$=\binom{n+1}{m}\left(\frac{m}{n+1}\right)^{m}\left(1-\frac{m}{n+1}\right)^{n+1-m}\,.$
The more interpretable upper bound on $\delta(n,m)$ is derived by substituting $m=c\cdot(n+1)$ in (3.2). The bound holds for $n\geq 1$ provided that $c(n+1)$ is an integer. It is obtained by upper bounding the error on Stirling’s approximation of the binomial coefficient, i.e.
$\binom{n+1}{c(n+1)}\leq\frac{e}{2\pi c^{c(n+1)}(1-c)^{(1-c)(n+1)}\sqrt{c(1-c)(n+1)}}\,.$
@@
$\mathbb{P}_{\mathcal{E}}\{Y_{n+1}\in\widehat{C}^{\widehat{\mathfrak{X}}(\mathcal{D}_{1:{n}})}_{n,m}(X_{n+1})|X_{n+1}\in\mathscr{X}_{k},{\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})}\}\leq 1-\alpha+\frac{1}{m-2}.$
@@
We first consider when $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ is used for the partition. Note as a result of Lemma 2 that $\{S_{i}\mid\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1}):i\in[n+1]\}$ are exchangeable random variables as the conditioning statistic $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ is a symmetric function of the data $(X_{1},S_{1}),\ldots,(X_{n+1},S_{n+1})$. Consider $\mathscr{X}_{k}\in\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$. Applying (A.5), we also have that $\{S_{i}\mid\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1}),X_{n+1}\in\mathscr{X}_{k}:i\in\mathscr{X}_{k}\}$, the conformity scores of data within the leaf node containing $X_{n+1}$, are conditionally exchangeable. We finally note that after conditioning on $\mathscr{X}_{k}$, the event $\mathcal{E}_{\mathrm{cand}}$ is symmetric in the indices $[n+1]$. Therefore, the random variables $\{S_{i}\mid\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1}),X_{n+1}\in\mathscr{X}_{k},\mathcal{E}_{\mathrm{cand}}:i\in\mathscr{X}_{k}\}$ are exchangeable as well. We denote the number of data points in the leaf $k$ as $m_{k}:=|\{X_{i}:X_{i}\in\mathscr{X}_{k}\}|$. This means that the rank of $S_{n+1}$ within $\mathscr{X}_{k}$ is uniformly distributed on the set $\{1,\ldots,m_{k}\}$. Explicitly, letting $\widetilde{R}_{n+1}$ denote the rank of $S_{n+1}$ within $\mathscr{X}_{k}$,
@@
The proof is by combining Lemma 1 with Lemma 5, by which we can still nearly enjoy the prescribed level of conditional coverage even if we are not conditioning on $\mathcal{E}$.
@@
When the conformity scores have a continuous joint distribution, the upper bound is shown analogously using the upper bound (A.15) in the proof of Theorem 3.1.
Appendix B Tighter Coverage Bounds with a Refitting Algorithm
We also consider a variation from Algorithm 2 that achieves tighter coverage bounds. The only difference is that for any $x_{n+1}\in\mathcal{X}$, the value of $\widehat{C}_{n,m}(x_{n+1})$ relies on a robust tree which has been fitted with the knowledge of the value $x_{n+1}$. The intuition behind including a value of $X_{n+1}$ is that it reduces a case in which the tree may potentially differ when $(X_{n+1},S_{n+1})$ is excluded. In particular, an additional candidate node can no longer be created by the number of points included increasing from $m-1$ to $m$, since the knowledge of $X_{n+1}$ has been incorporated.
B.1 Refitting Algorithm
@@
1:Fitted model $\hat{\mu}$, calibration dataset $\mathcal{D}_{1:n}$, tree hyperparameters ($m,K,\eta=0.05$), test point $X_{n+1}$
@@
6:Construct a prediction set $\widehat{C}_{n,m}$ as
$\widehat{C}_{n,m}(x)=\{y:S(X_{n+1},y)\leq S^{*}(x)\},$
@@
Algorithm 3 differs from Algorithm 1 in that it uses a value for $X_{n+1}$ by hallucinating a conformity score $\widetilde{S}_{n+1}$ that will never affect the criterion in any node in which $X_{n+1}$ lies. For example, $\widetilde{S}_{n+1}$ can be reset prior to each splitting decision to be equal to any other $S_{i}$ lying inside the same node as $X_{n+1}$. Because of this, $\widetilde{S}_{n+1}$ will never affect the criterion value in any node, and the tree will be fit as if the data point $(X_{n+1},\widetilde{S}_{n+1})$ were excluded from any criterion calculations.
This algorithm avoids any possibility of $(X_{n+1},S_{n+1})$ changing the tree by changing the candidacy status of a potential leaf node. The computational cost is equivalent when computing a prediction set for a single value of $X_{n+1}$, but scales linearly in the number of test points at which to evaluate $\widehat{C}_{n,m}(\cdot)$, whereas Algorithm 1 has constant cost as a function of the number of test points at which to evaluate $\widehat{C}_{n,m}(\cdot)$.
Lemma 6.
Let $(X_{1},Y_{1}),\ldots,(X_{n+1},Y_{n+1})$ be i.i.d. samples from a joint probability measure $\mathbb{P}$. Let $\widehat{\mathfrak{X}}_{n}(X_{n+1})$ denote the partition resulting from Algorithm 3 on data $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ and $X_{n+1}$. Let $\widehat{\mathfrak{X}}_{n+1}$ denote the partition resulting from Algorithm 2 on $(X_{1},Y_{1}),\ldots,(X_{n+1},Y_{n+1})$. Let $\mathscr{X}_{k(X_{n+1})}$ be the leaf of $\widehat{\mathfrak{X}}_{n}(X_{n+1})$ containing $X_{n+1}$. Define
@@
We first consider when $\widehat{\mathfrak{X}}_{n+1}$ is used as the partition. Note as a result of Lemma 2 that $\{S_{i}\mid\widehat{\mathfrak{X}}_{n+1}:i\in[n+1]\}$ are exchangeable random variables as the conditioning statistic $\widehat{\mathfrak{X}}_{n+1}$ is a symmetric function of the data $(X_{1},S_{1}),\ldots,(X_{n+1},S_{n+1})$. Applying (A.5), we have also that $\{S_{i}|\widehat{\mathfrak{X}}_{n+1},X_{n+1}\in\mathscr{X}_{k}:i\in\mathscr{X}_{k}\}$, the conformity scores of data within the leaf node containing $X_{n+1}$, are conditionally exchangeable. We denote the number of data points in leaf $k$ as $m_{k}:=|\{X_{i}:X_{i}\in\mathscr{X}_{k}\}|$. This means that the rank of $S_{n+1}$ within $\mathscr{X}_{k}$ is uniformly distributed on the set $\{1,\ldots,m_{k}\}$. Explicitly, letting $\widetilde{R}_{n+1}$ denote the rank of $S_{n+1}$ within $\mathscr{X}_{k}$,
@@
Now, by noting that on $\mathcal{E}_{\mathrm{inlier}}$ we have $\widehat{\mathfrak{X}}_{n+1}=\widehat{\mathfrak{X}}_{n}(X_{n+1})$, we conclude the proof. ∎
Theorem B.1.
Let $(X_{1},Y_{1}),\ldots,(X_{n+1},Y_{n+1})$ be i.i.d. samples from a joint probability measure $\mathbb{P}$. Assume we apply Algorithm 3 on $(X_{1},S(X_{1},Y_{1})),\ldots,(X_{n},S(X_{n},Y_{n}))$ for test point $X_{n+1}$ with resulting partition $\widehat{\mathfrak{X}}(X_{n+1})$. Then with $\widehat{C}_{n,m}^{\widehat{\mathfrak{X}}_{n}(X_{n+1})}(X_{n+1})$ obtained using Algorithm 3, we have
$\mathbb{P}\{Y_{n+1}\in\widehat{C}_{n,m}^{\widehat{\mathfrak{X}}_{n}(X_{n+1})}(X_{n+1})\mid X_{n+1}\in\mathscr{X}_{k},\widehat{\mathfrak{X}}(X_{n+1})\}\geq 1-\alpha-\frac{2}{m}\,.$
$\mathbb{P}\{Y_{n+1}\in\widehat{C}_{n,m}^{\widehat{\mathfrak{X}}_{n}(X_{n+1})}(X_{n+1})\mid X_{n+1}\in\mathscr{X}_{k},\widehat{\mathfrak{X}}(X_{n+1})\}\leq 1-\alpha+\frac{1}{m-2}+\frac{2}{m}\,.$
@@
We have
[…]
As for the upper bound, when the conformity scores have a continuous joint distribution, we have
[…]
Appendix C Experimental Details
C.1 Conformal Prediction for Regression

### v2

@@ 6 Concluding Remarks
Conformal prediction serves to provide some statistical clarity in the otherwise uncharted waters of modern black-box deep learning. Conformal tree allows for local adaptation and conditional coverage guarantees based on a self-grouping procedure that is robust to the inclusion of an additional i.i.d. test point. This can be a valuable tool for quantifying uncertainty in black-box learning applications where the model outputs a single point prediction using training data that has been obscured from the end user. Future work may include extending the framework beyond trees to other classes of calibration models.
@@ A.2 Proof of Lemma 1
Note also that a tree implementation with a minimum improvement threshold for splitting also requires condition 2 to be satisfied for the resulting tree to be different. This is because the binary variable of a split candidate’s criterion value being above the threshold changing necessarily requires the criterion value of the node to change as well. For the first case, the only possibility is that new candidates are created if the exclusion of $(X_{n+1},S_{n+1})$ changes the number of datapoints inside a potential leaf from $m$ to $m-1$. For this to occur, it requires that exactly $m$ points are contained in the potential leaf node containing $X_{n+1}$. Denote by $\mathscr{X}_{k(X_{n+1})}$ the leaf of $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ containing $X_{n+1}$. We define the event such that removing $X_{n+1}$ doest not change the candidacy (i.e. eligibility for splitting) of the node $\mathscr{X}_{k(X_{n+1})}$ as
@@
$\delta(n,m):=\frac{2}{m}+\exp\left\{-(2m-1)\left(1+\frac{1}{n}\right)\right\}$ (A.9)
@@
To upper bound the probability of $\mathcal{E}^{c}_{\mathrm{cand}}$, we consider an image similar to pigeonholes. Let the size of the partition be $K=|\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})|$. What we know is that, in $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$, each box should contain at least $m$ points. Therefore, we consider randomly assigning $n$ data points to the $K$ holes, while the $n+1$-th points will always go to the $k$-th hole. The event $\mathcal{E}^{c}_{\mathrm{cand}}$ is then the event that the $k$-th hole receives only $m-1$ points from the $n$ data points. We consider the following steps to scatter. Step 1) Choose $m*K-1$ balls and distribute them so that each hole should have exactly $m$ balls (counting the $n+1$-th point together). There are so many possible ways of doing it, and let’s denote each possible event as $A_{i}$ for $i=1,....,C$ for some $C$. Step 2) Distribute all remaining $n-Km+1$ balls except in the $k$ th hole, so that this hole remains exactly the size of $m$. Denote this event by $\mathcal{E}_{k}$. Then,
$\mathbb{P}(\mathcal{E}^{c}_{\mathrm{cand}})=\sum_{i=1}^{C}\mathbb{P}(A_{i})\mathbb{P}(\mathcal{E}_{k}|A_{i})=\sum_{i=1}^{C}\mathbb{P}(A_{i})\left(1-\frac{1}{K}\right)^{n-Km+1}=\left(1-\frac{1}{K}\right)^{n-Km+1}.$
By using $1-x\leq e^{-x}$ and $K\leq K_{\rm max}$, we have
$\left(1-\frac{1}{K}\right)^{n-Km+1}\leq e^{-\frac{n-Km+1}{K}}\leq\exp\left\{-\left(\frac{n+1}{K_{\rm max}}-m\right)\right\}.$
@@
$\mathbb{P}_{\mathcal{E}}\{Y_{n+1}\in\widehat{C}^{\widehat{\mathfrak{X}}(\mathcal{D}_{1:{n}})}_{n,m}(X_{n+1})|X_{n+1}\in\mathscr{X}_{k},{\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})}\}\leq 1-\alpha+\frac{1}{m+1}.$
@@
We first consider the properties of $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ before thinking of $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})$. The corresponding prediction interval in (2.1) with $\mathfrak{X}=\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ is denoted by $\widehat{C}^{\widehat{\mathfrak{X}}(\mathcal{D}_{1:{n+1}})}_{n,m}(X_{n+1})$. Note first as a result of Lemma 2 that $\{S_{i}\mid\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1}):i\in[n+1]\}$ are exchangeable random variables as the conditioning statistic $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ is a symmetric function of the data $(X_{1},S_{1}),\ldots,(X_{n+1},S_{n+1})$. Consider $\mathscr{X}_{k}\in\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$. Applying (A.5), we also have that $\{S_{i}\mid\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1}),X_{n+1}\in\mathscr{X}_{k}:i\in\mathscr{X}_{k}\}$, the conformity scores of data within the leaf node containing $X_{n+1}$, are conditionally exchangeable. We finally note that after conditioning on $\mathscr{X}_{k}$, the event $\mathcal{E}_{\mathrm{cand}}$ is symmetric in the indices $[n+1]$. Therefore, the random variables $\{S_{i}\mid\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1}),X_{n+1}\in\mathscr{X}_{k},\mathcal{E}_{\mathrm{cand}}:i\in\mathscr{X}_{k}\}$ are exchangeable as well. We denote the number of data points in the leaf $k$ as $m_{k}:=|\{X_{i}:X_{i}\in\mathscr{X}_{k}\}|$. This means that the rank of $S_{n+1}$ within $\mathscr{X}_{k}$ is uniformly distributed on the set $\{1,\ldots,m_{k}\}$. Explicitly, letting $\widetilde{R}_{n+1}$ denote the rank of $S_{n+1}$ within $\mathscr{X}_{k}$,
@@
The proof is by combining Lemma 1 with Lemma 5, by which we can still nearly enjoy the prescribed level of conditional coverage even if we are not conditioning on $\mathcal{E}$. Note that the notation $\widehat{C}_{n,m}(X_{n+1})$ in Theorem 3.1 is equivalent to $\widehat{C}^{\widehat{\mathfrak{X}}(\mathcal{D}_{1:{n}})}_{n.m}(X_{n+1})$ in Lemma 5.
@@
When the conformity scores have a continuous joint distribution, the upper bound is shown analogously using the upper bound in Theorem 3.1.
Appendix B Conformal Prediction through Robust Tree Refitting
B.1 Hallucination Algorithm
Here, we also consider a variation from Algorithm 2 to take $X_{n+1}$ into account without using the true $Y_{n+1}$ value. It leads to a tighter coverage bound, where the price is the computational cost for refitting the tree for every $X_{n+1}$ value.
@@ B.1 Hallucination Algorithm
1:Fitted model $\hat{\mu}$, calibration dataset $\mathcal{D}_{1:n}$, tree hyperparameters ($m,{K_{\rm max}},\eta=0.05$), test point $X_{n+1}$
@@
6:Construct a prediction set $\widehat{C}^{\prime}_{n,m}$ as
$\widehat{C}^{\prime}_{n,m}(x)=\{y:S(X_{n+1},y)\leq S^{*}(x)\},$
@@
Algorithm 3 differs from Algorithm 2 in that it uses a value for $X_{n+1}$ by hallucinating a conformity score $\widetilde{S}_{n+1}$ that will never affect the criterion in any node in which $X_{n+1}$ lies. For example, $\widetilde{S}_{n+1}$ can be reset prior to each splitting decision to be equal to any other $S_{i}$ lying inside the same node as $X_{n+1}$. Because of this, $\widetilde{S}_{n+1}$ will never affect the criterion value in any node, and the tree will be fit as if the data point $(X_{n+1},\widetilde{S}_{n+1})$ were excluded from any criterion calculations. Therefore, this algorithm avoids any possibility of $(X_{n+1},S_{n+1})$ changing the tree by changing the candidacy status of a potential leaf node. The computational cost is equivalent when computing a prediction set for a single value of $X_{n+1}$, but scales linearly in the number of test points at which to evaluate $\widehat{C}^{\prime}_{n,m}(\cdot)$, whereas Algorithm 1 has constant cost as a function of the number of test points at which to evaluate $\widehat{C}_{n,m}(\cdot)$.
B.2 Theoretical Coverage of Hallucination Algorithm
Now, we present that with the refitting algorithm in Algorithm 3, we have improved the bounds due to the additional knowledge on $X_{n+1}$. In this section, we use the following notations to denote partitions: Let $\widehat{\mathfrak{X}}_{n}(X_{n+1})=\widehat{\mathfrak{X}}(\mathcal{D}_{1:n},X_{n+1})$ denote the partition resulting from Algorithm 3 on data $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ and $X_{n+1}$. Let $\widehat{\mathfrak{X}}_{n+1}=\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$ denote the partition resulting from Algorithm 2 on $(X_{1},Y_{1}),\ldots,(X_{n+1},Y_{n+1})$.
Theorem B.1.
Under the same setting as in Theorem 3.1, but with the tree fitting algorithm as Algorithm 3 instead of Algorithm 2, denote the fitted partition as $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n},X_{n+1})=\{\mathscr{X}_{k}\}_{k=1}^{K}.$ Then with the prediction set $\widehat{C}^{\prime}_{n,m}(X_{n+1})$ defined in Algorithm 3, we have
$\mathbb{P}\{Y_{n+1}\in\widehat{C}^{\prime}_{n,m}(X_{n+1})\mid X_{n+1}\in\mathscr{X}_{k},\widehat{\mathfrak{X}}(\mathcal{D}_{1:n},X_{n+1})\}\geq 1-\alpha-\frac{2}{m}\,.$
$\mathbb{P}\{Y_{n+1}\in\widehat{C}^{\prime}_{n,m}(X_{n+1})\mid X_{n+1}\in\mathscr{X}_{k},\widehat{\mathfrak{X}}(\mathcal{D}_{1:n},X_{n+1})\}\leq 1-\alpha+\frac{1}{m-2}+\frac{2}{m}\,.$
@@ B.2 Theoretical Coverage of Hallucination Algorithm
We have
[…]
Now, we consider the upper bound. Note that on $\mathcal{E}_{\mathrm{inlier}}$ we have $\widehat{\mathfrak{X}}_{n+1}=\widehat{\mathfrak{X}}_{n}(X_{n+1})$ due to Lemma 7. Therefore, when the conformity scores have a continuous joint distribution, we have
[…]
@@
Note first as a result of Lemma 2 that $\{S_{i}\mid\widehat{\mathfrak{X}}_{n+1}:i\in[n+1]\}$ are exchangeable random variables as the conditioning statistic $\widehat{\mathfrak{X}}_{n+1}$ is a symmetric function of the data $(X_{1},S_{1}),\ldots,(X_{n+1},S_{n+1})$. Applying (A.5), we have also that $\{S_{i}|\widehat{\mathfrak{X}}_{n+1},X_{n+1}\in\mathscr{X}_{k}:i\in\mathscr{X}_{k}\}$, the conformity scores of data within the leaf node containing $X_{n+1}$, are conditionally exchangeable. We denote the number of data points in leaf $k$ as $m_{k}:=|\{X_{i}:X_{i}\in\mathscr{X}_{k}\}|$. This means that the rank of $S_{n+1}$ within $\mathscr{X}_{k}$ is uniformly distributed on the set $\{1,\ldots,m_{k}\}$. Explicitly, letting $\widetilde{R}_{n+1}$ denote the rank of $S_{n+1}$ within $\mathscr{X}_{k}$,
@@
Now, by Lemma 7, noting that on $\mathcal{E}_{\mathrm{inlier}}$ we have $\widehat{\mathfrak{X}}_{n+1}=\widehat{\mathfrak{X}}_{n}(X_{n+1})$, we conclude the proof. ∎
Lemma 7.
Let $(X_{1},Y_{1}),\ldots,(X_{n+1},Y_{n+1})$ be i.i.d. samples from a joint probability measure $\mathbb{P}$. Let $\mathscr{X}_{k(X_{n+1})}$ be the leaf of $\widehat{\mathfrak{X}}_{n}(X_{n+1})$ containing $X_{n+1}$. Define
@@
B.3 Empirical Consideration of Hallucination Algorithm
Split Conformal Conf. Tree (Hallucination) Conformal Tree Conf. Tree (CART) Conformal Forest Locally Weighted) CQR
Width Cov. Width Cov. P.B. Width Cov. P.B. Width Cov. P.B. Width Cov. P.B. Width Cov. P.B. Width Cov. P.B.
Data 1 4.06 0.9 3.24 0.86 0.66 3.24 0.86 0.66 4.04 0.93 0.54 3.85 0.9 0.54 4.31 0.91 0.55 3.82 0.89 0.64
Data 2 2.48 0.87 3.1 0.82 0.81 3.1 0.82 0.81 4.1 0.9 0.61 3.65 0.91 0.36 4.01 0.87 0.36 3.11 0.88 0.3
bike 2.05 0.93 2.0 0.9 0.67 2.06 0.91 0.66 2.4 0.95 0.43 2.12 0.93 0.49 2.17 0.95 0.46 1.9 0.91 0.57
bio 2.19 0.87 2.05 0.85 0.65 2.04 0.85 0.68 2.36 0.9 0.36 2.61 0.9 0.33 2.24 0.87 0.48 1.9 0.89 0.62
community 2.15 0.92 1.87 0.89 0.67 1.89 0.9 0.69 2.4 0.96 0.56 2.56 0.94 0.48 2.06 0.95 0.6 1.84 0.91 0.72
concrete 0.8 0.91 0.73 0.88 0.62 0.73 0.87 0.65 0.81 0.9 0.46 0.83 0.89 0.51 0.77 0.88 0.6 0.88 0.88 0.4
homes 0.9 0.9 0.93 0.88 0.79 0.96 0.88 0.77 1.28 0.9 0.64 1.17 0.9 0.46 0.84 0.9 0.69 0.93 0.91 0.53
star 0.18 0.9 0.16 0.85 0.7 0.16 0.86 0.7 0.2 0.93 0.24 0.2 0.91 0.44 0.2 0.92 0.35 0.2 0.92 0.28
Table 3: Comparison of width, coverage, and prediction bands (P.B.) across different methods and datasets, including the hallucination algorithm. The experiments were conducted on the randomly selected 500 data points in order to run the Hallucination algorithm in a reasonable amount of time.
We repeat the experiment described earlier in Section 4.2 using the hallucination version of the algorithm which boasts a tighter coverage guarantee. However, this algorithm is significantly more computationally expensive, as it requires the robust tree to be fit separately for each test point, rather than just once in total. For this reason, in this section, we limit the size of each dataset to 500 randomly chosen observations. The results are in Table 3.
We find the performance of Conformal Tree with and without hallucination to be very similar. This is because it will be rare that the inclusion of the covariate value for the test point adds a new eligible split and that this split is chosen among the split candidates. Both of these conditions are necessary for the tree from the hallucination algorithm to differ from the non-hallucination version. As this is fairly unlikely, we expect the two methods to perform similarly in most cases. While some of the intervals end up slightly sharper using the hallucination algorithm, we generally recommend against it if a low runtime is of concern to the practitioner. This is because the hallucination algorithm requires a new robust tree to be fit for each test point, which can add significant cost if the test set is large. On the other hand, if the user were only interested in a prediction set for a single test point, there would be almost no downside to the hallucination algorithm.
Appendix C Sensitivity Analysis for $K_{\rm max}$ and $m$
For each dataset, we vary $m$ and $K_{\rm max}$ over a grid of values in order to determine the relative effect of varying these hyperparameters on the performance of our method. We plot the resulting average interval score length on held-out data for each combination of $m$ and $K_{\rm max}$ in Figure 8. The figure shows that the best results are generally obtained with relatively small $m$, and often with the large $K_{\rm max}$ as possible, implying that there is a significant benefit in capturing more local structure in the data. With this in mind, we recommend that practitioners choose the smallest $m$ as possible where $m$ is chosen based on their desired worst-case coverage computed in Theorem 3.1. Practitioners are welcome to tune hyperparameters using additional held-out data if they so desire as well.
Figure 8: Comparison of varying values for $m$ and ${K_{\rm max}}$ on interval score length on each dataset. Metrics are computed on a held-out validation data split. The red-lined lined box indicates the optimal (achieving the smallest interval score length).
Appendix D Experimental Details
D.1 Details for Section 4.2